OpenStudy (anonymous):

Graph the inequality: -4x+3y<-12

OpenStudy (anonymous):

You need to put this in the slope intercept form to determine the line of your graph.

OpenStudy (anonymous):

i mean 12<-4x+3y

OpenStudy (anonymous):

Can you put in the slope / intercept form? y=mx+b

OpenStudy (anonymous):

y=mx+b ------> 12=-4x+3y

OpenStudy (anonymous):

Ok the formula is y= mx + b. So to convert the equation 4x+3y = -12you would first get y on the left side of the equation. So we would have 3y = -4x -12. Then we would divide both sides by 3 to get y= -4/3x -4. Using this formula you would then have to plot your line on the graph. Once it is posted, since it is an inequality you would have to shade the area above or below the line.

OpenStudy (anonymous):

thanks alot!